Corrections to “Reduced power....”

p.5 l.-2 after “...morphism” add a footnote saying “Another proof can be found in [?].”

p.7 l.-0 OK. p.8 l.9 after “this map is” add “a homomorphism which is” p.8 l.-3 replace “R” by “A” p.9 l.5 after “subcategory” add “of Sm/k which is” p.9 l.12 after “complement to” add “the closed subset” p.9 l.-12 before “such that” add “to zero” p.9 l.-10 after “+vt” add “)” p.9 l.-9 after “Consider the” add “flat” p.12 l.-15 add “is an isomorphism” before “from” p.12 l.-14 remove “is an isomorphism” p.13 l.14 put the equation after “Let” on a separate line p.13 l.-17 replace “O^{n}” with “O^{N}” p.13 l.-14 replace “is a cycle” by “cycle is” p.13 l.-13 replace “A_{X}^{n}” with “A_{X}^{N}” p.13 l.-11 replace “K_{n,R}” with “K_{N,r}” p.13 l.-6 replace “the fiber of” with “the fiber of the cycle representing” p.13 l.-6 replace “ã” with “a” p.13 l.-7 move everything starting with “Let” and ending by the period

on l.-5 to p.14 l.1 between the period and ’Indeed”

p.14 l.1 remove “actual” p.14 l.1 Replace “Indeed” by “Therefore” p.14 l.-8 introduce an equation number (label “neweq1” to be used below) p.15 In diagram (5.3) replace three vertical arrows by one long vertical

arrow p.16 l.10 after “subbundle” add “O → ξ” p.16 l.12 before “where” remove the subscript n from ξ p.16 l.13 remove everything from the beginning of the line to and including

“and”

p.16 l.14, l.-14, l.-13 (twice), l.-12, l.-10, l.-9, l.-7 remove the subscript n from ξ

p.16 l.-9 remove everything from “ξ_{n}” to and including “and” on the next line

p.16 l.-4 remove “The following lemma is straightforward.” p.17 l.4, l.5 remove the subscript n from ξ

p.17 l.6 Here and everywhere else below up to the end of the paper make μ bold

p.17 l.9 the dot at S should be higher S^{•}